# The Mass Distribution of Stars Stripped in Binaries: The Effect of Metallicity

### v1

@@ 4.4 Constraints from the Wolf-Rayet population
Adopting the mass thresholds for WR-type spectral morphology from Shenar et al. (2020, see also Table 1) and convolving with the adopted star formation rates, we estimate that there should be around , and (or , depending on whether the drop is present, see §4.3) WR stars produced from binary interaction in the Milky Way, LMC, and SMC, respectively.
The Milky Way has an observed sample of almost 333See the WR catalog by P.A. Crowther on https://pacrowther.staff.shef.ac.uk/WRcat/., which increases to after correcting for observational biases (Shenar, 2024). In contrast, the LMC and SMC WR samples are considered to be relatively complete, with 154 and 12 WR stars, respectively (Hainich et al., 2014, 2015; Shenar et al., 2016; Neugent et al., 2017; Shenar et al., 2019). Deshmukh et al. (2024) observed the binary fraction of WN-type WR stars in the Milky Way to be % (see also Dsilva et al., 2020, 2022, 2023), meaning that of the expected WR stars in the Milky Way should be WR binaries. The LMC has WR binary candidates (Shenar et al., 2019), and the SMC has (Shenar et al., 2016, 2018). These observed numbers of WR binaries match well with our predictions. However, the large fraction of the WR stars at all metallicities that cannot be explained by binary stripping further underscores the need to better understand the mechanisms behind WR star formation (Shenar et al., 2020).
@@ 4.5 Comparison with stripped star detections in the Magellanic Clouds
Drout et al. (2023) recently presented the results of an observational search for stripped stars in the Magellanic Clouds. Because candidates were identified through their excess UV emission compared to main-sequence stars, only systems with low extinction and lower-mass or dark companions could be identified. This means that the sample is incomplete, but it is still interesting to note that the most massive stripped star found in the SMC has (Götberg et al., 2023), which is just below the expected drop at . On the other hand, the WR stars in the SMC are substantially brighter and likely more massive (, Hainich et al., 2015; Shenar et al., 2016) than these stripped star candidates, leaving a gap in mass that appears empty of stripped stars. While higher mass stripped stars are more difficult to select via UV excess (due to their overlap with main sequence stars at higher luminosities), it is possible that this gap is the helium star desert (see §3.3), and that the WR stars formed through other pathways, such as enhanced mass loss via wind stripping or chemically-homogeneous evolution (§ 5.4). A better understanding of these mechanisms will be necessary to confirm the origins of the SMC’s WR stars.
A larger data set could help to confirm the existence of the helium star desert in other ways. With complete observations of stripped stars down to , it would be possible to constrain the slope of the stripped star mass distribution (see Fig. 4). While dependent on many factors, the slope would provide an additional constraint for uncertainties related to, for example, convective overshoot and the outcome of different envelope-stripping mechanisms. With an upcoming, newly-reduced UV photometry catalog (Ludwig et al., in prep.) based on data from the Swift SUMaC survey (Hagen et al., 2017), suitable data could soon be available. This will also be an interesting case for the UV mission UVEX, scheduled to launch in 2030, which will map the entire mass distribution of stripped stars (Kulkarni et al., 2021, see science pillar 1). However, we note that direct comparison to samples of hot stripped stars identified via UV excess will likely require further characterization of the expected contributions of stripped star plus compact object systems. While subdominant to the overall population (see §2.2), these systems are more efficiently identified via the UV excess method (Drout et al., 2023).
@@ 5 Implications and potential observational tests
5.1 Partially-stripped stars – the counterpart of hot stripped stars at low metallicity?
@@
In this study, we have considered the production of hot stripped stars through envelope-stripping in binaries that is initiated before central helium burning. However, at low metallicity, if a star is radiative and initiates mass transfer during or shortly before central helium burning, it could become a cool, partially-stripped star or continue mass transfer until stellar death (Klencki et al., 2022).
@@
These partially-stripped stars could be more numerous than hot stripped stars if they originate from high-mass donor stars and low-metallicity environments. However, this hinges on the prediction of late expansion for massive stars at low metallicity. As described in §4.1, a mass distribution with a drop predicts that the SMC will have only 5 hot binary-stripped stars with masses and initial masses , while the distribution without a drop predicts 20 such stars. This suggests that if the SMC does exhibit a drop, there should be stars with initial mass that entered nuclear timescale mass transfer or that are partially stripped. Since they likely would quickly lose a large fraction, but not the entirety, of their hydrogen-rich envelopes, we expect that their masses should be higher (e.g., ) than their hot counterparts (see, e.g., Fig. 11 of Klencki et al., 2022).
@@
Furthermore, these stars should be bright () and have temperatures that are similar to, or cooler than, the main sequence (kK). Because of their lower temperatures, partially-stripped stars should be significantly brighter in the optical wavelength range compared to hot stripped stars, which emit primarily in the UV (kK, Götberg et al., 2018). This suggests that partially-stripped stars are easily detectable but can masquerade as regular main-sequence stars or blue supergiants with somewhat lower surface gravity () and higher surface helium abundance () than expected (Klencki et al., 2022). However, while undergoing mass transfer, they could also be distinguished by appearing cooler than main-sequence stars and as eclipsing binaries and ellipsoidal variables (Prieto et al., 2008).
@@
In recent years, several intriguing observations have been presented of proposed partially-stripped stars in the SMC and LMC (Ramachandran et al., 2023, 2024). These partially-stripped stars appear as O- or B-type giants or supergiants and show surface enrichment in nitrogen and, in some cases, helium. An alternative explanation for these stars is that they are helium stars that have just undergone envelope-stripping and are observed while contracting before helium core burning starts. This evolutionary phase has been invoked to explain observations of several other stars (e.g., Irrgang et al., 2020; Bodensteiner et al., 2020; Irrgang et al., 2022; Villaseñor et al., 2023, see also Dutta & Klencki 2024). Mapping out the full parameter space of partially-stripped stars and such bloated stripped stars would help to constrain both late expansion and the outcome of binary interaction at low metallicity.
@@
It is worth noting as well that models predict that stars undergoing partial envelope-stripping or mass transfer that initiates only after the end of core helium burning should have a carbon-oxygen core structure similar to that of single stars and would be significantly different from that of stars stripped fully earlier in their evolution (Schneider et al., 2021; Klencki et al., 2022). Those core differences could affect these stars’ explodability and the formation of compact objects (Schneider et al., 2021; Vartanyan et al., 2021; Laplace et al., 2021, 2024).
@@ 5.2 Implications for high-redshift galaxies and the production of ionizing photons
Whether low-metallicity massive stars predominantly create hot stripped stars or cool partially-stripped stars directly impacts the output of ionizing radiation from stellar populations. This is because of the difference in effective temperature between these two binary products (cf. e.g., Götberg et al., 2023; Ramachandran et al., 2024). The ionizing output is of particular interest for high-redshift galaxies, both because ionizing radiation powered the reionization of the Universe, and because it is responsible for the nebular emission lines that are the main observable features of these distant populations. Metallicity is typically lower at high redshift than in most present-day galaxies, and most galaxies with redshift have a stellar metallicity at least as low as the SMC (Steidel et al., 2016; Strom et al., 2018; Cullen et al., 2019; Topping et al., 2020; Stanton et al., 2024). Therefore, the helium star desert, if real, should impact metal-poor, high-redshift galaxies.
@@
While we do not model the spectral energy distribution of massive stripped stars in this study, we use previous predicted ionizing emission rates to estimate the influence of the helium star desert on the hydrogen- and helium-ionizing output from a stellar population. We do this for a population with metallicity and constant star-formation rate of /yr, adopting the mass distributions presented in § 3.1.
@@
We use the ionizing emission rate estimates from Götberg et al. (2019), adopting photons per second as the total hydrogen-ionizing emission rate () from hot stripped stars with masses . In this population, the authors used the same assumptions for the binary population synthesis and the same underlying evolutionary models as we do here, but they only included stripped stars with masses up to 8.5 (corresponding to initial masses up to ). They interpolated over their spectral models to estimate the total ionizing emission from stripped stars across this mass range.
@@
To account for the contribution from higher mass stripped stars, we fit a straight line to as function of stripped star mass using the predictions for individual stripped stars with mass by Götberg et al. (2018). We then extrapolate this function out to to estimate the ionizing emission rates of the stripped stars with masses above 8.5 . We exclude stripped stars with mass , as these may have WR-type spectra (Shenar et al., 2020), which can have cooler photospheres located in the stellar wind and opaque stellar winds that block ionizing emission. This mass boundary is also below the mass where Sander & Vink (2020) estimate that wind mass loss rates significantly increase. Fig. 7 shows the model predictions and the extrapolation of the fit. We also display predictions for two other ionizing stars - an early O-type star and a hot WR star (Smith et al., 2002). The extrapolation falls below these numbers, but detailed evolutionary and spectral models would be necessary to provide a more accurate estimate.
@@
Using this extrapolation to assign ionizing emission rates to the stripped stars in our mass distribution, we predict that the stripped stars with should contribute hydrogen-ionizing photons per second in a population in which stars form at a constant rate of /yr. This constitutes a boost in the contribution from stripped stars compared to Götberg et al. (2019, see their Table A.3). If we instead use the mass distribution from the population to emulate a low- distribution without a drop in the mass distribution, we estimate that massive stripped stars would contribute hydrogen-ionizing photons per second for a constant star-formation rate of /yr. This would double the contribution from stripped stars compared to Götberg et al. (2019). However, in either case, stripped stars are not a dominant source of H-ionizing photon production from a full stellar population. Adopting the ionizing contribution from single stars from Starburst99 ( photons s-1, Leitherer et al. 1999), we estimate that the total contribution from stripped stars of all masses to the hydrogen-ionizing emission rate from a low- population should be between s-1 and s-1, or 6.5 to 9.1%, depending on whether or not the mass distribution exhibits a drop, respectively. These results are summarized visually in the top panel of Fig. 8.
@@
Stripped stars with transparent atmospheres, i.e., ones that do not show WR type spectra, could be expected to emit a substantial amount of He+-ionizing emission. We therefore make a rough estimate of the contribution of massive stripped stars to the helium-ionizing budget of a stellar population. We take the same approach as described previously in this section for the hydrogen-ionizing emission estimate (see Fig. 7). Again, we extrapolate up to , which we estimate has s-1. This agrees with the estimate for the hottest WN star (WR stars with nitrogen-dominated spectra) models from Smith et al. (2002) and the estimate for massive helium stars from Sander & Vink (2020). By summing the He+-ionizing emission for individual stars, we estimate the contribution from a population of massive stripped stars at to be photons per second for a constant star-formation rate of /yr. If, instead, late expansion does not occur at low metallicity, and the mass distribution does not have a drop (i.e., we adopt the mass distribution), the contribution could be as high as photons per second.
This exercise suggests that massive stripped stars without WR-type spectra could be the most prominent source of He+-ionizing emission from a stellar population. However, it is a rough estimate and observational confirmation would be required. The estimate for main sequence and WR stars from Starburst99 is photons s-1 and for stripped stars with it is photons s-1 (Götberg et al., 2019). Therefore, if there is no drop, the He+-ionizing budget could increase by 500%, reaching photons s-1. Even if there is a drop, the addition of massive stripped stars would still double the total budget of He+-ionizing photons, reaching a total of photons s-1 (Fig. 8, bottom panel). These values are similar to predictions from version 2.0 of the BPASS binary spectral population synthesis code (Stanway et al., 2016), included for comparison in Fig. 8. However, we note that the main ionizing source likely differs between our two estimates, since we do not include rapidly rotating, chemically homogeneously evolving accretor stars, while in BPASS they are prominent sources of ionizing emission.
@@ 5.3 Implications for the formation of gravitational wave progenitors
The progenitors of massive (20 M⊙) black holes are thought to originate in low metallicity environments (, Belczynski et al., 2010; Giacobbo & Mapelli, 2018; Klencki et al., 2018; Chruslinska et al., 2019; Neijssel et al., 2019; Broekgaarden et al., 2022). This suggests that these progenitors were either the massive hot stripped stars we model in this work or the partially-stripped or mass-transferring cool objects discussed in §5.1. While late expansion at low metallicity is included in rapid binary population synthesis codes (Pols et al., 1998; Hurley et al., 2000), the complex mass transfer phase associated with massive, low-metallicity stars merits more in-depth study and likely an updated treatment for population synthesis to ensure that the evolution in populations agrees with what is found in detailed evolutionary models (cf. e.g., Klencki et al., 2022; Ercolino et al., 2024). This could be particularly relevant to explore in context of the proposed evolutionary channel where stable mass transfer can bring two stars sufficiently close that a common envelope is not required for the future compact object system to merge within a Hubble time (van den Heuvel et al., 2017; Marchant et al., 2021).
@@
The late expansion of massive stars in binaries also determines whether mass transfer is initiated right after the end of the main sequence (i.e., Hertzprung gap donors) or only later during helium burning (i.e., core-He burning donors). Those two types of stars are commonly differentiated in rapid binary evolution codes (following Hurley et al., 2002). In population studies of gravitational wave sources, common envelope evolution has typically been limited to core-He burning donors (e.g., Dominik et al., 2012; Belczynski et al., 2016; Vigna-Gómez et al., 2020; Broekgaarden et al., 2021), with Hertzprung gap giants assumed to always lead to mergers and the formation of Thorne-Żytkow objects (cf. Farmer et al., 2023; O’Grady et al., 2024). As a result, the phenomenon of late expansion has an effect on the predicted formation rates of gravitational wave sources as a function of metallicity, contributing to the preference for the low metallicity. More recently, it has been demonstrated that the outcome of common envelope evolution is primarily determined by the type of the outer envelope (radiative vs. convective), rather than the evolutionary stage of the donor (Hertzprung gap vs. core-He burning) (Kruckow et al., 2016; Klencki et al., 2021; Romagnolo et al., 2024). It is still unclear, however, how the late expansion of massive stars affects the stable mass transfer channel. Klencki et al. (2022) found that mass transfer from a core-He burning donor is characterized by a lower mass exchange rate compared to that from a Hertzprung gap giant. This suggests an increased mass transfer stability (Klencki et al. in prep.), which is a crucial factor favoring the formation of close-orbit binary black hole systems that merge within a Hubble time (Marchant et al., 2021; Olejak et al., 2021).
5.4 Can rapidly rotating stars populate the helium star desert?

### v2

@@ 4.4 Constraints from the Wolf-Rayet population
Adopting the mass thresholds for WR-type spectral morphology from Shenar et al. (2020, see also Table 1) and convolving with the adopted star formation rates, we estimate that there should be around , and (or , depending on whether the drop is present, see Sect. 4.3) WR stars produced from binary interaction in the Milky Way, LMC, and SMC, respectively.
The Milky Way has an observed sample of almost 333See the WR catalog by P.A. Crowther on https://pacrowther.staff.shef.ac.uk/WRcat/., which increases to after correcting for observational biases (Shenar, 2024). In contrast, the LMC and SMC WR samples are considered to be relatively complete, with 154 and 12 WR stars, respectively (Hainich et al., 2014, 2015; Shenar et al., 2016; Neugent et al., 2017; Shenar et al., 2019). Deshmukh et al. (2024) observed the binary fraction of WN-type WR stars in the Milky Way to be % (see also Dsilva et al., 2020, 2022, 2023), meaning that of the expected WR stars in the Milky Way should be WR binaries and the remaining should be single. The LMC has WR binary candidates (Shenar et al., 2019), and the SMC has (Shenar et al., 2016, 2018). These observed numbers of WR binaries match well with our predictions. However, the large fraction of the WR stars at all metallicities that cannot be explained by binary stripping further underscores the need to better understand the mechanisms behind WR star formation (Shenar et al., 2020).
@@ 4.5 Comparison with stripped star detections in the Magellanic Clouds
Drout et al. (2023) recently presented the results of an observational search for stripped stars in the Magellanic Clouds. Because candidates were identified through their excess UV emission compared to main-sequence stars, only systems with low extinction and lower-mass or dark companions could be identified. This means that the sample is incomplete, but it is still interesting to note that the most massive stripped star found in the SMC has (Götberg et al., 2023), which is just below the expected drop at . On the other hand, the WR stars in the SMC are substantially brighter and likely more massive (, Hainich et al., 2015; Shenar et al., 2016) than these stripped star candidates, leaving a gap in mass that appears empty of stripped stars. While higher mass stripped stars are more difficult to select via UV excess (due to their overlap with main sequence stars at higher luminosities), it is possible that this gap is the helium star desert (see Sect. 3.3), and that the WR stars formed through other pathways, such as enhanced mass loss via wind stripping or chemically homogeneous evolution (Sect. 5.4). A better understanding of these mechanisms will be necessary to confirm the origins of the SMC’s WR stars.
A larger data set could help to confirm the existence of the helium star desert in other ways. With complete observations of stripped stars down to , it would be possible to constrain the slope of the stripped star mass distribution (see Fig. 4). While dependent on many factors, the slope would provide an additional constraint for uncertainties related to, for example, convective overshoot and the outcome of different envelope-stripping mechanisms. With an upcoming, newly reduced UV photometry catalog (Ludwig et al., in prep.) based on data from the Swift Ultraviolet Survey of the Magellanic Clouds (SUMaC, Hagen et al., 2017), suitable data could soon be available. This will also be an interesting case for the UV mission UVEX, scheduled to launch in 2030, which will map the entire mass distribution of stripped stars (Kulkarni et al., 2021, see science pillar 1). However, we note that direct comparison to samples of hot stripped stars identified via UV excess will likely require further characterization of the expected contributions of stripped star plus compact object systems. While subdominant to the overall population (see Sect. 2.2), these systems are more efficiently identified via the UV excess method (Drout et al., 2023).
@@ 5 Implications and potential observational tests
5.1 Partially stripped stars as the counterpart of hot stripped stars at low metallicity
@@
In this study, we have considered the production of hot stripped stars through envelope-stripping in binaries that is initiated before central helium burning. However, at low metallicity, if a star is radiative and initiates mass transfer during or shortly before central helium burning, it could become a cool, partially stripped star or continue mass transfer until stellar death (Klencki et al., 2022).
@@
These partially stripped stars could be more numerous than hot stripped stars if they originate from high-mass donor stars and low-metallicity environments. However, this hinges on the prediction of late expansion for massive stars at low metallicity. As described in Sect. 4.1, a mass distribution with a drop predicts that the SMC will have only 5 hot binary-stripped stars with masses and initial masses , while the distribution without a drop predicts 20 such stars. This suggests that if the SMC does exhibit a drop, there should be stars with initial mass that entered nuclear timescale mass transfer or that are partially stripped. Since they likely would quickly lose a large fraction, but not the entirety, of their hydrogen-rich envelopes, we expect that their masses should be higher (e.g., ) than their hot counterparts (see, e.g., Fig. 11 of Klencki et al., 2022).
@@
Furthermore, these stars should be bright () and have temperatures that are similar to, or cooler than, the main sequence (kK). Because of their lower temperatures, partially stripped stars should be significantly brighter in the optical wavelength range compared to hot stripped stars, which emit primarily in the UV (kK, Götberg et al., 2018). This suggests that partially stripped stars are easily detectable but can masquerade as regular main-sequence stars or blue supergiants with somewhat lower surface gravity () and higher surface helium abundance () than expected (Klencki et al., 2022). However, while undergoing mass transfer, they could also be distinguished by appearing cooler than main-sequence stars and as eclipsing binaries and ellipsoidal variables (Prieto et al., 2008).
@@
In recent years, several intriguing observations have been presented of proposed partially stripped stars in the SMC and LMC (Ramachandran et al., 2023, 2024). These partially stripped stars appear as O- or B-type giants or supergiants and show surface enrichment in nitrogen and, in some cases, helium. An alternative explanation for these stars is that they are helium stars that have just undergone envelope-stripping and are observed while contracting before helium core burning starts. This evolutionary phase has been invoked to explain observations of several other stars (e.g., Irrgang et al., 2020; Bodensteiner et al., 2020; Irrgang et al., 2022; Villaseñor et al., 2023, see also Dutta & Klencki 2024). Mapping out the full parameter space of partially stripped stars and such bloated stripped stars would help to constrain both late expansion and the outcome of binary interaction at low metallicity.
@@
It is worth noting as well that models predict that stars undergoing partial envelope-stripping or mass transfer that initiates only after the end of core helium burning should have a carbon-oxygen core structure similar to that of single stars and would be significantly different from that of stars stripped fully earlier in their evolution (Schneider et al., 2021; Klencki et al., 2022). Those core differences could affect these stars’ explodability and the formation of compact objects (Schneider et al., 2021; Vartanyan et al., 2021; Laplace et al., 2021, 2025).
@@ 5.2 Implications for high-redshift galaxies and the production of ionizing photons
Whether low-metallicity massive stars predominantly create hot stripped stars or cool partially stripped stars directly impacts the output of ionizing radiation from stellar populations. This is because of the difference in effective temperature between these two binary products (cf. e.g., Götberg et al., 2023; Ramachandran et al., 2024). The ionizing output is of particular interest for high-redshift galaxies, both because ionizing radiation powered the reionization of the Universe, and because it is responsible for the nebular emission lines that are the main observable features of these distant populations. Metallicity is typically lower at high redshift than in most present-day galaxies, and most galaxies with redshift have a stellar metallicity at least as low as the SMC (Steidel et al., 2016; Strom et al., 2018; Cullen et al., 2019; Topping et al., 2020; Stanton et al., 2024). Therefore, the helium star desert, if real, should impact metal-poor, high-redshift galaxies.
@@
While we did not model the spectral energy distribution of massive stripped stars in this study, we used previous predicted ionizing emission rates to estimate the influence of the helium star desert on the hydrogen- and helium-ionizing output from a stellar population. We did this for a population with metallicity and constant star formation rate of /yr, adopting the mass distributions presented in Sect. 3.1.
@@
We used the ionizing emission rate estimates from Götberg et al. (2019), adopting photons per second as the total hydrogen-ionizing emission rate () from hot stripped stars with masses . In this population, the authors used the same assumptions for the binary population synthesis and the same underlying evolutionary models as we did here, but they only included stripped stars with masses up to 8.5 (corresponding to initial masses up to ). They interpolated over their spectral models to estimate the total ionizing emission from stripped stars across this mass range.
@@
To account for the contribution from higher mass stripped stars, we fit a straight line to as function of stripped star mass using the predictions for individual stripped stars with mass by Götberg et al. (2018). We then extrapolated this function out to to estimate the ionizing emission rates of the stripped stars with masses above 8.5 . We excluded stripped stars with mass , as these may have WR-type spectra (Shenar et al., 2020), which can have cooler photospheres located in the stellar wind and opaque stellar winds that block ionizing emission. This mass boundary is also below the mass where Sander & Vink (2020) estimate that wind mass loss rates of stars significantly increase. Fig. 7 shows the model predictions and the extrapolation of the fit. We also display predictions for two other ionizing stars - an early O-type star and a hot WR star (Smith et al., 2002). The extrapolation falls below these numbers, but detailed evolutionary and spectral models would be necessary to provide a more accurate estimate.
@@
Using this extrapolation to assign ionizing emission rates to the stripped stars in our mass distribution, we predict that the stripped stars with should contribute hydrogen-ionizing photons per second in a population in which stars form at a constant rate of /yr. This constitutes a boost in the contribution from stripped stars compared to Götberg et al. (2019, see their Table A.3). If we instead use the mass distribution from the population to emulate a low- distribution without a drop in the mass distribution, we estimate that massive stripped stars would contribute hydrogen-ionizing photons per second for a constant star formation rate of /yr. This would double the contribution from stripped stars compared to Götberg et al. (2019). However, in either case, stripped stars are not a dominant source of H-ionizing photon production from a full stellar population. Adopting the ionizing contribution from single stars from Starburst99 ( photons s-1, Leitherer et al. 1999), we estimate that the total contribution from stripped stars of all masses to the hydrogen-ionizing emission rate from a low- population should be between s-1 and s-1, or 6.5 to 9.1%, depending on whether or not the mass distribution exhibits a drop, respectively. These results are summarized visually in the top panel of Fig. 8.
@@
Stripped stars with transparent atmospheres, that is, ones that do not show WR type spectra, could be expected to emit a substantial amount of He+-ionizing emission. We therefore made a rough estimate of the contribution of massive stripped stars to the helium-ionizing budget of a stellar population. We took the same approach as described previously in this section for the hydrogen-ionizing emission estimate (see Fig. 7). Again, we extrapolated up to , which we estimated has s-1. This agrees with the estimate for the hottest WN star (WR stars with nitrogen-dominated spectra) models from Smith et al. (2002) and the estimate for massive helium stars from Sander & Vink (2020). By summing the He+-ionizing emission for individual stars, we estimated the contribution from a population of massive stripped stars at to be photons per second for a constant star formation rate of /yr. If, instead, late expansion does not occur at low metallicity, and the mass distribution does not have a drop (i.e., we adopted the mass distribution), the contribution could be as high as photons per second.
This exercise suggests that massive stripped stars without WR-type spectra could be the most prominent source of He+-ionizing emission from a stellar population. However, it is a rough estimate and observational confirmation would be required. The estimate for main sequence and WR stars from Starburst99 is photons s-1 and for stripped stars with it is photons s-1 (Götberg et al., 2019). Therefore, if there is no drop, the He+-ionizing budget could increase by 500%, reaching photons s-1. Even if there is a drop, the addition of massive stripped stars would still double the total budget of He+-ionizing photons, reaching a total of photons s-1 (Fig. 8, bottom panel). These values are similar to predictions from version 2.0 of the BPASS binary spectral population synthesis code (Stanway et al., 2016), included for comparison in Fig. 8. However, we note that the main ionizing source likely differs between our two estimates, since we did not include rapidly rotating, chemically homogeneously evolving accretor stars, while in BPASS they are prominent sources of ionizing emission.
@@ 5.3 Implications for the formation of gravitational wave progenitors
The progenitors of massive (20 M⊙) black holes are thought to originate in low metallicity environments (, Belczynski et al., 2010; Giacobbo & Mapelli, 2018; Klencki et al., 2018; Chruslinska et al., 2019; Neijssel et al., 2019; Broekgaarden et al., 2022). This suggests that these progenitors were either the massive hot stripped stars we modeled in this work or the partially stripped or mass-transferring cool objects discussed in Sect. 5.1. While late expansion at low metallicity is included in rapid binary population synthesis codes (Pols et al., 1998; Hurley et al., 2000), the complex mass transfer phase associated with massive, low-metallicity stars merits more in-depth study and likely an updated treatment for population synthesis to ensure that the evolution in populations agrees with what is found in detailed evolutionary models (cf. e.g., Klencki et al., 2022; Ercolino et al., 2024). This could be particularly relevant to explore in context of the proposed evolutionary channel where stable mass transfer can bring two stars sufficiently close that a common envelope is not required for the future compact object system to merge within a Hubble time (Inayoshi et al. 2017; van den Heuvel et al. 2017; Marchant et al. 2021, see also Bavera et al. 2021).
@@
The late expansion of massive stars in binaries also determines whether mass transfer is initiated right after the end of the main sequence (i.e., Hertzprung gap donors) or only later during helium burning (i.e., core-He burning donors). Those two types of stars are commonly differentiated in rapid binary evolution codes (following Hurley et al., 2002). In population studies of gravitational wave sources, common envelope evolution has typically been limited to core-He burning donors (e.g., Dominik et al., 2012; Belczynski et al., 2016; Vigna-Gómez et al., 2020; Broekgaarden et al., 2021), with Hertzprung gap giants assumed to always lead to mergers and the formation of Thorne-Żytkow objects (cf. Farmer et al., 2023; O’Grady et al., 2024). As a result, the phenomenon of late expansion has an effect on the predicted formation rates of gravitational wave sources as a function of metallicity, contributing to the preference for low metallicity (though in the current models it is not the leading-order effect, see e.g., van Son et al., 2025). More recently, it has been demonstrated that the outcome of common envelope evolution is primarily determined by the type of the outer envelope (radiative vs. convective), rather than the evolutionary stage of the donor (Hertzprung gap vs. core-He burning) (Kruckow et al., 2016; Klencki et al., 2021; Romagnolo et al., 2025). It is still unclear, however, how the late expansion of massive stars affects the stable mass transfer channel. Klencki et al. (2022) found that mass transfer from a core-He burning donor is characterized by a lower mass exchange rate compared to that from a Hertzprung gap giant. This suggests an increased mass transfer stability (Klencki et al. in prep.), which is a crucial factor favoring the formation of close-orbit binary black hole systems that merge within a Hubble time (Marchant et al., 2021; Olejak et al., 2021; van Son et al., 2022; Picco et al., 2024).
5.4 Rapidly rotating stars and the helium star desert